# Generative Adversarial Networks for Image Super-Resolution: A Survey

### v4

@@ 3.1.2. GANs on big samples for object detection
Object detection has wide applications in the industry, i.e., smart transportation (Wang et al., 2021a) and medical diagnosis (Gerdprasert and Mabu, 2021), etc. However, complex environments have huge challenges for pursuing good performance of object detection methods (Zou et al., 2019). Rich data is important for object detection. Existing methods used a data-driven strategy to collect a large-scale dataset including different object examples under different conditions to obtain an object detector. However, the obtained dataset does not contain all kinds of deformed and occluded objects, which limits effects of object detection methods. To resolve the issue, GANs are used for object detection (Ehsani et al., 2018; Li et al., 2017a). Ehsani et al. used segmentation and generation in a GANs from invisible parts in the objects to overcome occluded objects (Ehsani et al., 2018). To address a challenge of small object detection on low-resolution and noisy representation, a perceptual GAN (Perceptual GAN) reduced differences of small objects and big objects to improve performance in small object detection (Li et al., 2017a). That is, its generator converted poor perceived representation from small objects to high-resolution big objects to fool a discriminator, where mentioned big objects are similar to real big objects (Li et al., 2017a). To obtain sufficient information of objects, an end-to-end multi-task generative adversarial network (SOD-MTGAN) used a generator to recover detailed information for generating high-quality images for achieving accurate detection (Bai et al., 2018). Also, a discriminator transferred classification and regression losses in a back-propagated way into a generator (Bai et al., 2018). Two operations can extract objects from backgrounds to achieve good performance in object detection [59]. More detailed information is shown in Table 3.
@@ 4. GANs for image super-resolutions
According to mentioned illustrations, it is clear that GANs have many important applications in image processing. Also, image super-resolution is crucial for high-level vision tasks, i.e., medical image diagnosis and weather forecast, etc. Thus, GANs in image super-resolution have important significance in the real world. However, there are few summaries about GANs for image super-resolution. Inspired by that, we show GANs for image super-resolution, according to supervised GANs, semi-supervised GANs and unsupervised GANs for image super-resolution as shown in Fig. 8. Specifically, supervised GANs in image super-resolution include supervised GANs based improved architectures, supervised GANs based prior knowledge, supervised GANs with improved loss functions and supervised GANs based multi-tasks for image super-resolution. Semi-supervised GANs for image super-resolution contain semi-supervised GANs based improved architectures, semi-supervised GANs with improved loss functions and semi-supervised GANs based multi-tasks for image super-resolution.
Unsupervised GANs for image super-resolution consists of unsupervised GANs based improved architectures, unsupervised GANs based prior knowledge, unsupervised GANs with improved loss functions and unsupervised GANs based multi-tasks in image super-resolution. More information of GANs on image super-resolution can be illustrated as follows.
@@ 4.1. Supervised GANs for image super-resolution
4.1.1. Supervised GANs based improved architectures for image super-resolution
@@
In terms of designing novel and discriminators and generators, progressive growing generative adversarial networks (PGGAN or ProGAN) utilized different convolutional layers to progressively enlarge low-resolution images to improve image qualities for image recognition (Karras et al., 2017). An enhanced SRGAN (ESRGAN) used residual dense blocks into a generator without batch normalization to mine more detailed information for image super-resolution (Wang et al., 2018b). To eliminate effects of checkerboard artifacts and the unpleasing high-frequency, multi-discriminators were proposed for image super-resolution (Lee et al., 2019). That is, a perspective discriminator was used to overcome checkerboard artifacts and a gradient perspective was utilized to address unpleasing high-frequency question in image super-resolution. To improve the perceptual quality of predicted images, ESRGAN+ fused two adjacent layers in a residual learning way based on residual dense blocks in a generator to enhance memory abilities and added noise in a generator to obtain stochastic variation and obtain more details of high-resolution images (Rakotonirina and Rasoanaivo, 2020).
Restoring detailed information may generate artifacts, which can seriously affect qualities of restored images (Zhang and Ling, 2020). In terms of face image super-resolution, Zhang et al. used a supervised pixel-wise GAN (SPGAN) to obtain higher-quality face images via given low-resolution face images of multiple scale factors to remove artifacts in image super-resolution (Zhang and Ling, 2020). In terms of remote sensing image super-resolution, Gong et al. used enlighten blocks to make a deep network achieve a reliable point and used self-supervised hierarchical perceptual loss to overcome effects of artifacts in remote sensing image super-resolution (Gong et al., 2021). Dharejo et al. used Wavelet Transform (WT) characteristics into a transferred GAN to eliminate artifacts to improve quality of predicted remote sensing images (Dharejo et al., 2021). Moustafa et al. embedded squeeze-and-excitation blocks and residual blocks into a generator to obtain more high-frequency details (Moustafa and Sayed, 2021). Besides, Wasserstein distance is used to enhance the stability of training a remote sensing super-resolution model (Moustafa and Sayed, 2021). To address pseudo-textures problem, a saliency analysis is fused with a GAN to obtain a salient map that can be used to distinguish difference between a discriminator and a generator (Ma et al., 2019).
@@
Figure 8. Frame of GANs for image super-resolution.
[…]
The structure of the Section 4.
@@
CMRI-CGAN (Xia et al., 2021a) CGAN CGAN with optical flow component for magnetic resonance image super-resolution
@@ 4.1.2. Supervised GANs based prior knowledge for image super-resolution
It is known that combination of discriminative method and optimization can make a tradeoff between efficiency and performance (Zhang et al., 2017). Guan et al. used high-resolution image to low-resolution image network and low-resolution image to high-resolution image network with nearest neighbor down-sampling method to learn detailed information and noise prior for image super-resolution (Guan et al., 2019). Chan et al. used rich and diverse priors in a given pretrained to mine latent representative information for generating realistic textures for image super-resolution (Chan et al., 2021b). Liu et al. used a gradient prior into a GAN to suppress the effect of blur kernel estimation for image super-resolution (Liu et al., 2019b).
@@ 4.1.4. Supervised GANs based multi-tasks for image super-resolution
Improving image quality is important for high-level vision tasks, i.e., image recognition (Talab et al., 2019). Besides, devices often suffer from effects of multiple factors, i.e., device hardware, camera shakes and shooting distances, which results in collected images are damaged. That may include noise and low-resolution pixels. Thus, addressing the multi-tasks for GANs are very necessary (Hu et al., 2019). For instance, Adil et al. exploited SRGAN and a denoising module to obtain a clear image. Then, they used a network to learn unique representative information for identifying a person (Adil et al., 2020). In terms of image super-resolution and object detection, Wang et al. used multi-class cyclic super-resolution GAN to restore high-quality images, and used a YOLOv5 detector to finish object detection task (Wang et al., 2022). Zhang et al. used a fully connected network to implement a generator for obtaining high-definition plate images and a multi-task discriminator is used to enhance super-resolution and recognition tasks (Zhang et al., 2018b). The use of an adversarial learning was a good tool to simultaneously address text recognition and super-resolution (Xu et al., 2020).
In terms of complex damaged image restoration, GANs are good choices (Li et al., 2022). For instance, Li et al. used a multi-scale residual block and an attention mechanism in a GAN to remove noise and restore detailed information in CTA image super-resolution (Li et al., 2022). Nneji et al. improved a VGG19 to fine-tune two sub-networks with a wavelet technique to simultaneously address COVID-19 image denoising and super-resolution problems (Nneji et al., 2022). More information is shown in Table 7.
@@
JPLSRGAN (Zhang et al., 2018b) GAN GAN for license plate recognition and image super-resolution
@@
MRD-GAN (Li et al., 2022) GAN GAN with attention mechanism for image super-resolution and denoising
@@
RealESRGAN (Wang et al., 2021a) ESRGAN ESRGAN with pure synthetic data for blind image super-resolution
@@ 4.2. Semi-supervised GANs for image super-resolution
4.2.1. Semi-supervised GANs based improved architectures for image super-resolution
For real problems with less data, semi-supervised techniques are developed. For instance, asking patients takes multiple CT scans with additional radiation doses to conduct paired CT images for training SR models in clinical practice is not realistic. Motivated by that, GANs in semi-supervised ways are used for image super-resolution (You et al., 2019a). For instance, by maintaining the cycle-consistency of Wasserstein distance, a mapping from noisy low-resolution images to high-resolution images was built (You et al., 2019a). Besides, combining a convolutional neural network, residual learning operations in a GAN can facilitate more detailed information for image super-resolution (You et al., 2019a). To resolve super-resolution and few labeled samples, Xia et al. used soft multi-labels to implement a semi-supervised super-resolution method for person re-identification (Xia et al., 2021b). That is, first, a GAN is used to conduct a SR model. Second, a graph convolutional network is exploited to construct relationship of local features from a person. Third, some labeled samples are used to train unlabeled samples via a graph convolutional network.
@@ 4.2.1. Semi-supervised GANs based improved architectures for image super-resolution
In terms of multi-tasks, Nicolo et al. proposed to use a mixed adversarial Gaussian domain adaptation in a GAN in a semi-supervised way to obtain more useful information for implementing a 3D super-resolution and segmentation (Savioli et al., 2021). More information of semi-supervised GANs in image super-resolution can be illustrated in Table 8.
Table 8. Semi-supervised GANs for image super-resolution in section 4.2.
@@
MSSR (Xia et al., 2021b) GAN GAN with soft multi-labels in a semi-supervised way for image super-resolution
@@ 4.3. Unsupervised GANs for image super-resolution
Collected images in the real world have less pairs. To address this phenomenon, unsupervised GANs are presented (Yuan et al., 2018). It can be divided into four types, i.e., improved architectures, prior knowledge, loss functions and multi-tasks in GANs in unsupervised ways for image super-resolution as follows.
@@
4.3.1. Unsupervised GANs based improved architectures for image super-resolution
@@
CycleGANs have obtained success in unsupervised ways in image-to-image translation applications (Zhu et al., 2017). Accordingly, the CycleGANs are extended into SISR to address unpair images (i.e., low-resolution and high-resolution) in the datasets in the real world (Yuan et al., 2018). Yuan et al. used a CycleGAN for blind super-resolution over the following phases (Yuan et al., 2018). The first phase removed noise from noisy and low-resolution images. The second phase resorted to an up-sampled operation in a pre-trained deep network to enhance the obtained low-resolution images. The third phase used a fine-tune mechanism for a GAN to obtain high-resolution images. To address blind super-resolution, bidirectional structural consistency was used into a GAN in an unsupervised way to train a blind SR model and construct high-quality images (Zhao et al., 2018). Alternatively, Zhang et al. exploited multiple GANs as basis components to implement an improved CycleGAN for train an unsupervised SR model (Zhang et al., 2019c). To eliminate checkerboard artifacts, an upsampling module containing a bilinear interpolation and a transposed convolution was used in an unsupervised CycleGAN to improve visual effects of restored images in the real world (Kim et al., 2020).
@@
There are also other popular methods that use GANs in unsupervised ways for image super-resolution (Prajapati et al., 2020). To improve the learning ability of a SR model in the real world, it combines an unsupervised learning and a mean opinion score in a GAN to improve perceptual quality in the real-world image super-resolution (Prajapati et al., 2020). To recover more natural image characteristics, Lugmayr et al. combined unsupervised and supervised ways for blind image super-resolution (Lugmayr et al., 2019). The first step learned to invert the effects of bicubic down sampling operation in a GAN in an unsupervised way to extract useful information from natural images (Lugmayr et al., 2019). To generate image pairs in the real world, the second step used a pixel-wise network in a supervised way to obtain high-resolution images (Lugmayr et al., 2019). To break fixed downscaling kernel, Sefi et al. used KernelGAN (Bell-Kligler et al., 2019) and Internal-GAN (Shocher et al., 2018) to obtain an internal distribution of patches in the blind image super-resolution. To accelerate the training speed, a guidance module was used in a GAN to quickly seek a correct mapping from a low-resolution domain to a high-resolution domain in unpaired image super-resolution (Lian et al., 2019). To improve the accuracy of medical diagnosis, Song et al. used dual GANs in a self-supervised way to mine high dimensional information for PET image super-resolution (Mahapatra et al., 2017). Besides, other SR methods can have an important reference value for unsupervised GANs with for image super-resolution. For example, Wang et al. used an unsupervised method to translate real low-resolution images to real low-resolution images (Wang et al., 2020b). Chen et al. resorted to a supervised super-resolution method to convert obtained real low-resolution images into real high-resolution images (Chen et al., 2020). More information of mentioned unsupervised GANs for image super-resolution can be shown in Table 9 as follows.
@@
Table 9. Unsupervised GANs based improved architectures for image super-resolution.
@@
TrGAN (Wang et al., 2020b) GAN Unsupervised GAN for image synthesis and super-resolution

### v5

@@ 3.1.2. GANs on big samples for object detection
Object detection has wide applications in the industry, i.e., smart transportation (Wang et al., 2021b) and medical diagnosis (Gerdprasert and Mabu, 2021), etc. However, complex environments have huge challenges for pursuing good performance of object detection methods (Zou et al., 2019). Rich data is important for object detection. Existing methods used a data-driven strategy to collect a large-scale dataset including different object examples under different conditions to obtain an object detector. However, the obtained dataset does not contain all kinds of deformed and occluded objects, which limits effects of object detection methods. To resolve the issue, GANs are used for object detection (Ehsani et al., 2018; Li et al., 2017a). Ehsani et al. used segmentation and generation in a GANs from invisible parts in the objects to overcome occluded objects (Ehsani et al., 2018). To address a challenge of small object detection on low-resolution and noisy representation, a perceptual GAN (Perceptual GAN) reduced differences of small objects and big objects to improve performance in small object detection (Li et al., 2017a). That is, its generator converted poor perceived representation from small objects to high-resolution big objects to fool a discriminator, where mentioned big objects are similar to real big objects (Li et al., 2017a). To obtain sufficient information of objects, an end-to-end multi-task generative adversarial network (SOD-MTGAN) used a generator to recover detailed information for generating high-quality images for achieving accurate detection (Bai et al., 2018). Also, a discriminator transferred classification and regression losses in a back-propagated way into a generator (Bai et al., 2018). Two operations can extract objects from backgrounds to achieve good performance in object detection. More detailed information is shown in Table 3.
@@ 4. GANs for image super-resolutions
According to mentioned illustrations, it is clear that GANs have many important applications in image processing. Also, image super-resolution is crucial for high-level vision tasks, i.e., medical image diagnosis and weather forecast, etc. Thus, GANs in image super-resolution have important significance in the real world. However, there are few summaries about GANs for image super-resolution. Inspired by that, we show GANs for image super-resolution, according to supervised GANs, semi-supervised GANs and unsupervised GANs for image super-resolution as shown in Fig 2. Specifically, supervised GANs in image super-resolution include supervised GANs based designed network architectures, supervised GANs based prior knowledge, supervised GANs with improved loss functions and supervised GANs based multi-tasks for image super-resolution. Semi-supervised GANs for image super-resolution contain semi-supervised GANs based designed network architectures, semi-supervised GANs with improved loss functions and semi-supervised GANs based multi-tasks for image super-resolution.
Unsupervised GANs for image super-resolution consists of unsupervised GANs based designed network architectures, unsupervised GANs based prior knowledge, unsupervised GANs with improved loss functions and unsupervised GANs based multi-tasks in image super-resolution. More information of GANs on image super-resolution can be illustrated as follows.
@@ 4.1. Supervised GANs for image super-resolution
4.1.1. Supervised GANs based designed network architectures for image super-resolution
@@
In terms of designing novel and discriminators and generators, progressive growing generative adversarial networks (PGGAN or ProGAN) utilized different convolutional layers to progressively enlarge low-resolution images to improve image qualities for image recognition (Karras et al., 2017). To achieve better visual quality with more realistic and natural textures, an enhanced SRGAN (ESRGAN) used residual dense blocks into a generator without batch normalization to mine more detailed information for image super-resolution (Wang et al., 2018b). To eliminate effects of checkerboard artifacts and the unpleasing high-frequency, multi-discriminators were proposed for image super-resolution (Lee et al., 2019). That is, a perspective discriminator was used to overcome checkerboard artifacts and a gradient perspective was utilized to address unpleasing high-frequency question in image super-resolution. To improve the perceptual quality of predicted images, ESRGAN+ fused two adjacent layers in a residual learning way based on residual dense blocks in a generator to enhance memory abilities and added noise in a generator to obtain stochastic variation and obtain more details of high-resolution images (Rakotonirina and Rasoanaivo, 2020).
Figure 8. Frame of GANs for image super-resolution.
Restoring detailed information may generate artifacts, which can seriously affect qualities of restored images (Zhang and Ling, 2020). The methods mentioned in this paragraph can effectively alleviate this phenomenon. In terms of face image super-resolution, Zhang et al. used a supervised pixel-wise GAN (SPGAN) to obtain higher-quality face images via given low-resolution face images of multiple scale factors to remove artifacts in image super-resolution (Zhang and Ling, 2020). In terms of remote sensing image super-resolution, Gong et al. used enlighten blocks to make a deep network achieve a reliable point and used self-supervised hierarchical perceptual loss to overcome effects of artifacts in remote sensing image super-resolution (Gong et al., 2021). Dharejo et al. used Wavelet Transform (WT) characteristics into a transferred GAN to eliminate artifacts to improve quality of predicted remote sensing images (Dharejo et al., 2021). Moustafa et al. embedded squeeze-and-excitation blocks and residual blocks into a generator to obtain more high-frequency details (Moustafa and Sayed, 2021). Besides, Wasserstein distance is used to enhance the stability of training a remote sensing super-resolution model (Moustafa and Sayed, 2021). To address pseudo-textures problem, a saliency analysis is fused with a GAN to obtain a salient map that can be used to distinguish difference between a discriminator and a generator (Ma et al., 2019).
@@
CMRI-CGAN (Xia et al., 2021b) CGAN CGAN with optical flow component for magnetic resonance image super-resolution
@@ 4.1.2. Supervised GANs based prior knowledge for image super-resolution
It is known that combination of discriminative method and optimization can make a tradeoff between efficiency and performance (Zhang et al., 2017). To handle the blind distortions in real low resolution images, Guan et al. used high-resolution image to low-resolution image network and low-resolution image to high-resolution image network with nearest neighbor down-sampling method to learn detailed information and noise prior for image super-resolution (Guan et al., 2019). Chan et al. used rich and diverse priors in a given pretrained to mine latent representative information for generating realistic textures for image super-resolution (Chan et al., 2021b). Liu et al. used a gradient prior into a GAN to suppress the effect of blur kernel estimation for image super-resolution (Liu et al., 2019b).
@@ 4.1.4. Supervised GANs based multi-tasks for image super-resolution
Improving image quality is important for high-level vision tasks, i.e., image recognition (Talab et al., 2019). Besides, devices often suffer from effects of multiple factors, i.e., device hardware, camera shakes and shooting distances, which results in collected images are damaged. That may include noise and low-resolution pixels. Thus, addressing the multi-tasks for GANs are very necessary (Hu et al., 2019). For instance, Adil et al. exploited SRGAN and a denoising module to obtain a clear image. Then, they used a network to learn unique representative information for identifying a person (Adil et al., 2020). In terms of image super-resolution and object detection, Wang et al. used multi-class cyclic super-resolution GAN to restore high-quality images, and used a YOLOv5 detector to finish object detection task (Wang et al., 2022). Zhang et al. used a fully connected network to implement a generator for obtaining high-definition plate images and a multi-task discriminator is used to enhance super-resolution and recognition tasks (Zhang et al., 2018a). The use of an adversarial learning was a good tool to simultaneously address text recognition and super-resolution (Xu et al., 2020).
In terms of complex damaged image restoration, GANs are good choices (Li et al., 2022b). For instance, Li et al. used a multi-scale residual block and an attention mechanism in a GAN to remove noise and restore detailed information in CTA image super-resolution (Li et al., 2022b). Nneji et al. improved a VGG19 to fine-tune two sub-networks with a wavelet technique to simultaneously address COVID-19 image denoising and super-resolution problems (Nneji et al., 2022). More information is shown in Table 7.
@@
JPLSRGAN (Zhang et al., 2018a) GAN GAN for license plate recognition and image super-resolution
@@
MRD-GAN (Li et al., 2022b) GAN GAN with attention mechanism for image super-resolution and denoising
@@
RealESRGAN (Wang et al., 2021b) ESRGAN ESRGAN with pure synthetic data for blind image super-resolution
@@ 4.2. Semi-supervised GANs for image super-resolution
4.2.1. Semi-supervised GANs based designed network architectures for image super-resolution
For real problems with less data, semi-supervised techniques are developed. For instance, asking patients takes multiple CT scans with additional radiation doses to conduct paired CT images for training SR models in clinical practice is not realistic. Motivated by that, GANs in semi-supervised ways are used for image super-resolution (You et al., 2019a). For instance, by maintaining the cycle-consistency of Wasserstein distance, a mapping from noisy low-resolution images to high-resolution images was built (You et al., 2019a). Besides, combining a convolutional neural network, residual learning operations in a GAN can facilitate more detailed information for image super-resolution (You et al., 2019a). To resolve super-resolution with few labeled samples, Xia et al. used soft multi-labels to implement a semi-supervised super-resolution method for person re-identification (Xia et al., 2021a). That is, first, a GAN is used to conduct a SR model. Second, a graph convolutional network is exploited to construct relationship of local features from a person. Third, some labeled samples are used to train unlabeled samples via a graph convolutional network.
@@ 4.2.1. Semi-supervised GANs based designed network architectures for image super-resolution
In terms of multi-tasks, Nicolo et al. proposed to use a mixed adversarial Gaussian domain adaptation in a GAN in a semi-supervised way to obtain more useful information for implementing a 3D super-resolution and segmentation (Savioli et al., 2021). This method achieved better performance on many metrics. More information of semi-supervised GANs in image super-resolution can be illustrated in Table 8.
Table 8. Semi-supervised GANs for image super-resolution in section $2.$
@@
MSSR (Xia et al., 2021a) GAN GAN with soft multi-labels in a semi-supervised way for image super-resolution
@@ 4.3. Unsupervised GANs for image super-resolution
Collected images in the real world have less pairs, which suffers from challenge for supervised GANs in SISR. To address this phenomenon, unsupervised GANs are presented (Yuan et al., 2018). Unsupervised GANs for image super-resolution can be divided into six types: improved architectures, prior knowledge, loss functions, multi-task learning, real image super-resolution and real-world LR image synthesis in GAN-based unsupervised super-resolution, as follows.
@@
4.3.1. Unsupervised GANs based designed network architectures for image super-resolution
@@
CycleGANs have obtained success in unsupervised ways in image-to-image translation applications (Zhu et al., 2017). Accordingly, the CycleGANs are extended into SISR to address unpair images (i.e., low-resolution and high-resolution) in the datasets in real world (Yuan et al., 2018). Yuan et al. used a CycleGAN for blind super-resolution over the following phases (Yuan et al., 2018). The first phase removed noise from noisy and low-resolution images. The second phase resorted to an up-sampled operation in a pre-trained deep network to enhance the obtained low-resolution images. The third phase used a fine-tune mechanism for a GAN to obtain high-resolution images. To address blind super-resolution, bidirectional structural consistency was used into a GAN in an unsupervised way to train a blind SR model and construct high-quality images (Zhao et al., 2018). Alternatively, Zhang et al. exploited multiple GANs as basis components to implement an improved CycleGAN for train an unsupervised SR model (Zhang et al., 2019c). And achieves comparable performance with many state-of-the-art supervised models.
@@
There are also other popular methods that use GANs in unsupervised ways for image super-resolution (Prajapati et al., 2020). To improve the learning ability of a SR model in the real world, it combines an unsupervised learning and a mean opinion score in a GAN to improve perceptual quality in the real world image super-resolution (Prajapati et al., 2020). To break fixed downscaling kernel, KernelGAN (Bell-Kligler et al., 2019), TVG-KernelGAN(Park et al., 2023a), and Internal-GAN (Shocher et al., 2018) are used to obtain an internal distribution of patches in the blind image super-resolution. To accelerate the training speed, a guidance module was used in a GAN to quickly seek a correct mapping from a low-resolution domain to a high-resolution domain in unpaired image super-resolution (Lian et al., 2019). To improve the accuracy of medical diagnosis, Song et al. used dual GANs in a self-supervised way to mine high dimensional information for PET image super-resolution (Mahapatra et al., 2017). Besides, other SR methods can have an important reference value for unsupervised GANs with for image super-resolution. For example, to improve both image synthesis quality and representation learning performance under the unsupervised setting, Wang et al. used an unsupervised method to translate real low-resolution images to real low-resolution images (Wang et al., 2020a). To achieve better degradation learning and super-resolution performance, Chen et al. resorted to a supervised super-resolution method to convert obtained real low-resolution images into real high-resolution images (Chen et al., 2020). More information of mentioned unsupervised GANs for image super-resolution can be shown in Table 9 as follows.
@@
Table 9. Unsupervised GANs based designed network architectures for image super-resolution.
@@
TrGAN (Wang et al., 2020a) GAN Unsupervised GAN for image synthesis and super-resolution